OpenStudy (3psilon):

Help with math of this? Find T , 2T *sin45 = Mg

OpenStudy (3psilon):

@Algebraic!

OpenStudy (anonymous):

wut?

OpenStudy (anonymous):

\[T= \frac{ Mg }{ 2 \sin 45 }\]

OpenStudy (3psilon):

The answer says it = Mg/2 what happens to the root2 / 2

OpenStudy (anonymous):

no idea what you're asking... tbh

OpenStudy (3psilon):

hahha Just wanna know how they got from \[2T *\frac{ \sqrt{2} }{ 2 } = Mg \] to\[T = \]

OpenStudy (3psilon):

mg/2

OpenStudy (3psilon):

idk whats happening to the root 2 over 2

OpenStudy (3psilon):

Yeah , but I don't see how they're getting rid of the square root?

OpenStudy (anonymous):

\[\frac{ \sqrt{2}Mg }{ 2 }\]

OpenStudy (anonymous):

right?

OpenStudy (3psilon):

yes

OpenStudy (anonymous):

so, I guess you're asking:\[\sqrt{2}T =Mg\] \[T = \frac{Mg }{ \sqrt{2}}\] \[T = \frac{Mg }{ \sqrt{2}} \frac{ \sqrt{2} }{ \sqrt{2} }\]

OpenStudy (3psilon):

But they got T = Mg/2 in the book that I'm using?

OpenStudy (3psilon):

I guess it's just a typo

OpenStudy (anonymous):

possibly.

OpenStudy (anonymous):

maybe the answer is actually Mg/2 ... I don't know.

OpenStudy (3psilon):

Hahah thanks I know the questions was very vague but thanks :)

OpenStudy (anonymous):

did you want me to look the actual problem?

OpenStudy (3psilon):

I'll draw it, or attempt to

OpenStudy (3psilon):